The IQ test results of a group of students in a certain college were taken and are presented in a frequency distribution below.


Classes f

70-75 2

76-81 8

82-87 19

88-93 21

94-99 28

100-105 38

106-111 15

112-117 9


1. Use the frequency distribution above, determine the following:

a. Size of the class interval

b. Lower boundary of the median class

C. Median class

d. Number of students with an IQ greater than 99.5


2. Compute the value of:

A. Mean

B. Median

C. Mode


1
Expert's answer
2021-12-23T12:11:13-0500

1.

a. The class width for a class in a frequency distribution is found by subtracting the lower (or upper) class limit of one class from the lower (or upper) class limit of the next class


"76-70=6"

b.

Calculate the total frequency


"n=2+8+19+21+28+38+15+9=140"

Calculate "n\/2"


"n\/2=140\/2=70"

Calculate cumulative frequency. Select Cumulative frequency just greater than or equal to "n\/2."


"2+8+19+21=50,"




"2+8+19+21+28=78,"

We find that the 70th observation lies in the class "94-99."

Lower boundary of the median class is "l=93.5."


c. Median class is "93.5-99.5."


d.


"38+15+9=62"

There are 62 students with an IQ greater than 99.5.


2.

A.


"\bar{x}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nf_ix_i"

"=\dfrac{1}{140}(2(72.5)+8(78.5)+19(84.5)+21(90.5)"

"+28(96.5)+38(102.5)+15(108.5)+9(114.5))"

"=\dfrac{13534}{140}\approx96.6714"

B.

"l=" lower boundary point of median class "=93.5"

"n=" Total frequency "=140"

"cf=" Cumulative frequency of the class preceding the median class "=50"

"f=" Frequency of the median class "=28"

"c=" class length of median class "=6"


"Median\ M=l+\dfrac{n/2-cf}{f}\cdot c"

"=93.5+\dfrac{70-50}{28}\cdot 6\approx97.7857"

C.

To find Mode Class

Here, maximum frequency is 38.

The mode class is "99.5-105.5."

"L=" lower boundary point of mode class "=99.5"

"f_1="  frequency of the mode class "=38"

"f_0=" frequency of the preceding class "=28"

"f_2="  frequency of the succedding class "=15"

"c=" class length of mode class "=6"


"Z=L+(\dfrac{f_1-f_0}{2f_1-f_0-f_2})\cdot c"

"=99.5+(\dfrac{38-28}{2(38)-28-15})\cdot 6c"

"\approx101.3182"